OpenStudy (anonymous):

I really need help understanding how to do this. I will give medal and become a fan if you help me. :)

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\large 8^{x-2}=3\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

for this you need two things 1) the change of base formula which says \[\large b^x=y\iff x=\frac{\log(y)}{\log(b)}\] and 2) a calculator

OpenStudy (anonymous):

\[\large 8^{2-x}=3\iff 2-x=\frac{\log(3)}{\log(8)}\] solve for \(x\) and using algebra, then use a calculator to get the numeric answer

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

So I would do log(3)/log(8) which gets .53 - 2 that gets -1.47 and then multiply it be -1 to get 1.47?

OpenStudy (anonymous):

yes i guess so i would do the algebra first, but either way should do it

OpenStudy (anonymous):

\[x=2-\frac{\log(3)}{\log(8)}\] which is what you got too

OpenStudy (anonymous):

Okay thank you so much

OpenStudy (anonymous):

yw